the complement of an angle is six less than twice the measure of an angle itself. find the angle and its complement.
Answer by BurningFlame_marc(77)

Hello!
I will just show you how to solve this.
Representation: let x=angle, 2x-6= the complement.
Equation: x+(2x-6)=90
Solution: x+2x-6=90
3x-6=90
3x=96
x=32
Final Answer: The angle is 32 and the complement is 58.